# The Marcinkiewicz-Zygmund Property for Riemann Differences with Geometric Nodes

### v1

@@
Abstract.
We study when a Riemann difference of order $n$ possesses the Marcinkiewicz-Zygmund (MZ) property: that is, whether the conditions $f(h)=o(h^{n-1})$ and $Df(h)=o(h^{n})$ imply $f(h)=o(h^{n})$. This implication is known to hold for some classical examples with geometric nodes, such as $\{0,1,q,\dots,q^{n-1}\}$ and $\{1,q,\dots,q^{n}\}$, leading to a conjecture that these are the only such Riemann differences with the MZ property (see [2]). However, this conjecture was disproved by the third-order example with nodes $\{-1,0,1,2\}$ ([5]), and we provide further counterexamples and a general classification here.
We establish a complete analytic criterion for the MZ property by developing a recurrence framework: we analyze when a function $R(h)$ satisfying $D(h)=R(qh)-AR(h)$, together with $D(h)=o(h^{n})$ and $R(h)=o(h^{n-1})$, forces $R(h)=o(h^{n})$. We prove that this holds if and only if $A$ lies outside a critical modulus annulus determined by $q$ and $n$, covering both $|q|>1$ and $|q|<1$ cases. This leads to a complete characterization of all Riemann differences with geometric nodes that possess the MZ property, and provides a flexible analytic framework applicable to broader classes of generalized differences.
Key words and phrases:
Peano derivative, Riemann difference, MZ property, Taylor expansion
2020 Mathematics Subject Classification:
Primary 26A24; Secondary 26A27
1. Introduction
@@
(1.1) $\sum_{k=1}^{n+1}a_{k}b_{k}^{j}=\begin{cases}0&\text{for }j=0,1,\dots,n-1,\\ 1&\text{for }j=n.\end{cases}$
@@
which exists at $c=0$ for $f(x)=|x|$ but $f$ fails to be differentiable at 00.
@@
follows from Taylor’s theorem and the structure of (1.1).
@@
$f_{k}(c)\text{ exists for }0\leq k\leq n-1\text{ and }Rf^{(n)}(c)\text{ exists}\Rightarrow f_{n}(c)\text{ exists}.$
@@
$D(h)=R(qh)-AR(h),$
and determine precisely when such a relation, combined with decay conditions $D(h)=o(h^{n})$ and $R(h)=o(h^{n-1})$, forces $R(h)=o(h^{n})$. Our main result shows that this happens if and only if $|A|$ lies outside a critical annulus determined by $q$ and $n$. The classification applies uniformly across the cases $|q|>1$ and $10<|q|<$ 1, and leads to a necessary and sufficient condition for the MZ property in terms of the roots of the characteristic polynomial of the difference.
@@ 2. Necessary conditions
where $d(x),d_{i}\in F$ and $h_{i}\in H$. Define $f(x)=d(x)$, the coefficient of $\alpha$ in the Hamel expansion of $x$. Then $f$ satisfies (2.1), since for any $x,y\in\mathbb{R}$,
@@
Because $f(\alpha)=1$ and $f(\beta)=0$, $f$ is not linear (indeed $f(0)=0$), and hence $f$ is nowhere continuous and non-measurable.
@@
by the system (1.1). Hence $Rf^{(n)}(c)=0$ for all $c\in\mathbb{R}$.
This function shows that the implication $Rf^{(n)}(c)\Rightarrow f_{n}(c)$ can fail everywhere when $f$ is not continuous at $c$. Thus, for the implication to hold, it is necessary that $f$ be continuous at $c$. More generally, as the next example illustrates, the existence of the first $n-1$ Peano derivatives at $c$ is also necessary.
Example 2.2.
Let $n\geq 2$, let $\{b_{1},\ldots,b_{n+1}\}$ be a set of nodes, and let $F$ be the field from Example 2.1. Define
@@
Then $f$ is measurable and satisfies $f(c+h)=o(h^{n-2})$ as $h\to 0$. By the conditions (1.1), we again have $Rf^{(n)}(c)=0$. However, $f_{n-1}(c)=(n-1)!$ on $F$ and 00 on its complement, so $f_{n}(c)$ fails to exist.
We now formalize the obstruction to the MZ property. Suppose there exists a function $f$ and a point $c$ such that $f_{k}(c)$ exists for $k=0,\dots,n-1$, and $Rf^{(n)}(c)$ exists, but $f_{n}(c)$ does not. Define a function
@@
Then $g_{k}(0)=0$ for all $k<n$, and $R^{(n)}g(0)=0$ by the system (1.1). Moreover, $g_{n}(0)$ exists if and only if $f_{n}(c)$ exists. Thus, if $R$ is defined by
$Rg(h)=\sum_{k}a_{k}g(b_{k}h),$
then we may rephrase the MZ property as follows:
The difference $R$ has the Marcinkiewicz-Zygmund (MZ) property if, for all functions $g$, the conditions $g(h)=o(h^{n-1})$ and $Rg(h)=o(h^{n})$ imply $g(h)=o(h^{n})$.
Notation. In what follows, we distinguish between two related uses of the symbol $R$. When we write $Rf(h)$, we mean the application of a linear difference operator $R$, defined by
@@
where $\{a_{k},b_{k}\}$ are fixed coefficients and nodes. If these coefficients satisfy the moment system (1.1), then $R$ is called a Riemann difference of order $n$.
@@
In contrast, when we write $R(h)$, we refer to an abstract complex-valued function of $h$, which may or may not arise from applying a difference operator to a function. Theorems involving recursions like $D(h)=R(qh)-AR(h)$ operate at this more general level, where both $R(h)$ and $D(h)$ are understood as functions satisfying asymptotic and recurrence properties. This distinction allows us to formulate results in maximum generality while still applying them to concrete difference operators.
@@
Lemma 2.3.
Let $R$ be a difference of order $n$ with coefficients $a_{k}$ and nodes $b_{k}$, and let $c\neq 0$, $r\neq 0$. Define the rescaled difference $R^{\prime}$ by
$R^{\prime}f(h):=\sum_{k}ra_{k}f\left(\frac{b_{k}}{c}h\right).$
@@
Convention. Throughout this paper, we write $R(h)=o(h^{n})$ to mean that
$\lim_{h\to 0}\frac{|R(h)|}{|h|^{n}}=0,$
where $R(h)$ is complex-valued and defined on $\mathbb{R}\setminus\{0\}$. Similarly, if $Rf(h)=\sum a_{k}f(b_{k}h)$ is a difference operator applied to a function $f$, we say $Rf(h)=o(h^{n})$ if the same asymptotic relation holds.
3. Recurrence criteria for asymptotic decay
@@
  • In the intermediate range $|q|^{n-1}<|A|\leq|q|^{n}$, we construct explicit counterexamples where $R(h)=o(h^{n-1})$ and $D(h)=o(h^{n})$ but $R(h)\not=o(h^{n})$.
@@
under which the recurrence $D(h)=R(qh)-AR(h)$ is transformed into the equivalent relation
@@
3.1. Decay when $A$ lies outside the critical annulus
Theorem 3.1 (Low-growth case).

### v2

@@
Abstract
We study when a Riemann difference of order $n$ possesses the
Marcinkiewicz-Zygmund (MZ) property: that is, whether the conditions $f(h)=o(h^{n-1})$ and $Df(h)=o(h^{n})$ imply $f(h)=o(h^{n})$. This implication is known to hold for some classical examples with geometric nodes, such as $\{0,1,q,\dots,q^{n-1}\}$ and $\{1,q,\dots,q^{n}\}$, leading to a conjecture that these are the only such Riemann differences with the MZ property (see [2]). However, this conjecture was disproved by the third-order example with nodes $\{-1,0,1,2\}$ ([5]), and we provide further counterexamples and a general classification here.
We establish a complete analytic criterion for the MZ property by developing a recurrence framework: we analyze when a function $R(h)$ satisfying D(h)=R(qh)AR(h)D(h)=R(qh)-AR(h), together with $D(h)=o(h^{n})$ and $R(h)=o(h^{n-1})$, forces $R(h)=o(h^{n})$. We prove that this holds if and only if $A$ lies outside a critical modulus annulus determined by $q$ and $n$, covering both $|q|>1$ and $|q|<1$ cases. This leads to a complete characterization of all Riemann differences with geometric nodes that possess the MZ property, and provides a flexible analytic framework applicable to broader classes of generalized differences.
keywords:
Peano derivative
keywords:
Riemann difference
keywords:
MZ property
1 Introduction
@@
$\sum_{k=1}^{n+1}a_{k}b_{k}^{j}=\begin{cases}0&\text{for }j=0,1,\dots,n-1,\\ 1&\text{for }j=n.\end{cases}$ (1)
@@
which exists at $c=0$ for $f(x)=|x|$ but $f$ fails to be differentiable at 0.
@@
follows from Taylor’s theorem and the structure of (1).
@@
[…]
@@
D(h)=R(qh)AR(h),D(h)=R(qh)-AR(h),
and determine precisely when such a relation, combined with decay conditions $D(h)=o(h^{n})$ and $R(h)=o(h^{n-1})$, forces $R(h)=o(h^{n})$. Our main result shows that this happens if and only if $|A|$ lies outside a critical annulus determined by $q$ and $n$. The classification applies uniformly across the cases $|q|>1$ and $0<|q|<1$, and leads to a necessary and sufficient condition for the MZ property in terms of the roots of the characteristic polynomial of the difference.
@@ 2 Necessary conditions
where $d(x),d_{i}\in F$ and $h_{i}\in H$. Define $f(x)=d(x)$, the coefficient of $\alpha$ in the Hamel expansion of $x$. Then $f$ satisfies (2), since for any $x,y\in\mathbb{R}$,
@@
Because $f(\alpha)=1$ and $(\beta)=0, ff$ is not linear (indeed $f(0)=0$), and hence $f$ is nowhere continuous and non-measurable.
@@
by the system (1). Hence $Rf^{(n)}(c)=0$ for all $c\in\mathbb{R}$.
This function shows that the implication $Rf^{(n)}(c)\Rightarrow f_{n}(c)$ can fail everywhere when $f$ is not continuous at $c$. Thus, for the implication to hold, it is necessary that $f$ be continuous at $c$. More generally, as the next example illustrates, the existence of the first n1n-1 Peano derivatives at $c$ is also necessary.
Example 2.
Let $n\geq 2$, let $\{b_{1},\ldots,b_{n+1}\}$ be a set of nodes, and let $F$ be the field from Example 1. Define
@@
Then $f$ is measurable and satisfies $f(c+h)=o(h^{n-2})$ as $h\to 0$. By the conditions (1), we again have $Rf^{(n)}(c)=0$. However, $f_{n-1}(c)=(n-1)!$ on $F$ and 0 on its complement, so $f_{n}(c)$ fails to exist.
Suppose there exists a function $f$ and a point $c$ such that $f_{k}(c)$ exists for $k=0,\dots,n-1$, and $Rf^{(n)}(c)$ exists. Define a function
@@
Then $g_{k}(0)=0$ for all $k<n$, and $R^{(n)}g(0)=0$ by the system (1). This is equivalent to the asymptotic conditions
$g(h)=o(h^{n-1})\text{ and }Rg(h)=o(h^{n})$
where $Rg(h)=\sum_{k}a_{k}g(b_{k}h)$. Moreover, $f_{n}(c)$ exists if and only if $g_{n}(0)$ exists in which case $g(h)=o(h^{n})$. This observation motivates the following general definition of the Marcinkiewicz-Zygmund property.
Definition 3.
Let $\{a_{k}\}_{k=1}^{m}$ be a finite sequence of complex numbers, and let $\{b_{k}\}_{k=1}^{m}$ be a finite sequence of distinct real numbers. The associated linear difference operator $R$ is defined by
$Rg(h)=\sum_{k}a_{k}g(b_{k}h)$
where the $a_{k}$ are called coefficients, and the $b_{k}$ are the nodes of the operator.
Let $n\geq 1$. We say that $R$ has the Marcinkiewicz–Zygmund (MZ) property of order $n$ if, for all real-valued functions $g$ the conditions
$g(h)=o(h^{n-1})\text{ and }Rg(h)=o(h^{n})$
imply
$g(h)=o(h^{n}).$
Here, $f(h)=o(h^{n})$ means $\lim_{h\to 0}\frac{|f(h)|}{|h|^{n}}=0.$
Notation and Clarification. In what follows, we fix a positive integer $n$, so references to the MZ property will omit the explicit mention of order $n$, which is understood from context.
We distinguish between two related uses of the symbol $R$. When we write $Rf(h)$, we mean the application of a linear difference operator $R$, defined by
@@
where $\{a_{k},b_{k}\}$ are fixed coefficients and nodes. If these coefficients satisfy the moment system (1), then $R$ is called a Riemann difference of order $n$.
@@
In contrast, when we write $R(h)$, we refer to an abstract complex-valued function of $h$, which may or may not arise from applying a difference operator to a function. Theorems involving recursions like D(h)=R(qh)AR(h)D(h)=R(qh)-AR(h) operate at this more general level, where both $R(h)$ and $D(h)$ are understood as functions satisfying asymptotic and recurrence properties. This distinction allows us to formulate results in maximum generality while still applying them to concrete difference operators.
@@
Lemma 4.
Let $R$ be a difference with coefficients $a_{k}$ and nodes $b_{k}$, and let $c\neq 0$, $r\neq 0$. Define the rescaled difference $R^{\prime}$ by
$R^{\prime}f(h):=\sum_{k}ra_{k}f\left(cb_{k}h\right)=rRf(ch).$
@@
3 Recurrence criteria for asymptotic decay
@@
  • In the intermediate range $|q|^{n-1}<|A|\leq|q|^{n}$, ($|q|^{n}\leq|A|<|q|^{n}$ if $|q|<1$) we construct explicit counterexamples where $R(h)=o(h^{n-1})$ and $D(h)=o(h^{n})$ but $R(h)\not=o(h^{n})$.
@@
under which the recurrence D(h)=R(qh)AR(h)D(h)=R(qh)-AR(h) is transformed into the equivalent relation
@@
3.1 Decay when $A$ lies outside the critical annulus
Theorem 5 (Low-growth case).